OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=9, (4,1)

OpenStudy (anonymous):

y-y1=m(x-x1) plug in points for x1 and y1 and slope for m

OpenStudy (anonymous):

so how would that look?

OpenStudy (anonymous):

where there is an m, put a 9. Where there is an y1 put a 1 where there is an x1 put a 4.

myininaya (myininaya):

replace m with 9 replace x1 with 4 replace y1 with 1

OpenStudy (anonymous):

I don't understand what is so magical about searching and replacing that gives people such a hard time.

OpenStudy (anonymous):

9x-4(1)?

myininaya (myininaya):

you need to call on the powers of merlin

OpenStudy (anonymous):

The point they give is (x1,y1) = (4,1) Therefore x1 is 4 y1 is 1 Therefore y - y1 = m(x-x1) IS y - 1 = 9(x-4)

OpenStudy (anonymous):

Just find and replace